An air craft flight at an attitude of 320m it fire a rocket with the initial velocity of 300km/h how long does it take the rocket to reach the ground

1
Expert's answer
2022-05-24T00:23:01-0400
"L=v_0\sqrt{2H/g}\\
=(300\div3.6\rm\: m/s)\sqrt{2*320\:m/9.8\: m/s^2}=673\: m"